A 26.3-kg child jumps off a diving board 1.5m above the surface of a pool. If all of the child’s potential energy is converted into kinetic energy, how fast is the child going when he hits the water?


1
Expert's answer
2022-01-19T12:40:16-0500
"\dfrac 12mv^2=mgh,\\\space\\
v=\sqrt{2gh}=\sqrt{2·9.8·1.5}=5.42\text{ m/s}."